OpenStudy (anonymous):

Prove that sqrt(3) is an irrational number

OpenStudy (anonymous):

Assume 3 = a^2/b^2 (a, b in lowest terms) and try to get a contradiction.

OpenStudy (anonymous):

If I'm doing this right, I get to a^2/b^2=3b^2/b^2, conclude that because this can be simplified the fraction a/b is not in its lowest terms. But this would work the same if you used 4 instead of 3.

OpenStudy (mayankdevnani):

ok @dr25161

OpenStudy (mayankdevnani):

fine @dr25161

OpenStudy (anonymous):

thanks. that works

OpenStudy (mayankdevnani):

ok welcome

OpenStudy (anonymous):

3 = a^2/b^2 -> 3b^2 = a^2 So both a^2 and b^2 are divisible by 3 and not in lowest terms.